# Integrated Multimodal Sensing and Communication: Challenges, Technologies, and Architectures

### v1

@@ V Case Study
To assess the effectiveness of the examined schemes, we conduct a case study showcasing vision-RF multimodal fusion for enhancing ISAC performance. To support efficient multimodal fusion and centralized semantic reasoning, the F-MAC scheme is adopted. As illustrated in Fig. 3(a), we consider a surveillance scenario involving a visual sensor (camera), an RF sensor (radar), a base station, and multiple sensing targets. Each sensor is equipped with a lightweight edge agent that extracts high-level features from its respective data modalities (e.g., images and radar echoes), which are subsequently transmitted to the base station via SC. At the base station, a LAM functions as the central agent, performing multimodal feature fusion and semantic-level decoding to identify the targets and estimate their motion parameters, including distance, velocity, and azimuth angle. Based on these sensing outcomes, the base station adaptively steers the ISAC beam toward the targets, thereby enhancing the overall communication rate.
Figure 3: F-MAC-based case study. (a) The illustration of the implementation scenario. (b) Key parameter settings.
@@ V-A Experimental Settings
To construct a specialized dataset for training and evaluation, we use the VIRAT Video Dataset [11], which provides real-world surveillance video in diverse outdoor scenes. We first select three representative video clips, sampling one frame per second, resulting in approximately 10,000 RGB frames. We assume that all sensing targets in the scenes are vehicles.
To extract training labels, we first apply the YOLOv10 model [12] to identify bounding boxes of all vehicles in each frame. Next, using the Segment Anything Model (SAM) [13], we extract precise foreground vehicle segments based on those bounding boxes, generating around 800,000 labeled vehicle images for image reconstruction tasks. In addition to visual data, radar-related annotations are synthesized. We assume a fixed radar location at the lower-right corner of each image. For each vehicle, we calculate its distance and azimuth based on the relative position in the frame and assign it a velocity profile. Simulated RF signals are then generated accordingly. Some key parameter settings are shown in Fig. 3(b).
We compare it against a single-modality baseline, RF-ISAC, a conventional unimodal approach that relies exclusively on RF signals for sensing. The evaluation employs root mean squared error (RMSE) as the metric for sensing accuracy and communication rate as the metric for communication performance.
@@ V-B Evaluation Results
Figure 4: RMSE Comparison Between F-MAC and RF-ISAC Across SNRs.
Figure 5: The communication rate comparisons between F-MAC and RF-ISAC across different SNRs.
Fig. 4 presents a quantitative comparison of the RMSE performance for motion parameter estimation between F-MAC and RF-ISAC across varying SNR levels. The results demonstrate that F-MAC consistently outperforms RF-ISAC across all three key metrics: azimuth, distance, and velocity. Specifically, F-MAC maintains remarkably stable and low RMSE values, with azimuth errors around $2.2\times 10^{-3}$, distance errors near $1.3\times 10^{-3}$, and velocity errors close to $1.2\times 10^{-3}$, while exhibiting minimal fluctuation across different SNRs. In contrast, RF-ISAC shows higher RMSE, particularly under the velocity estimation. On average and across all SNR levels, F-MAC achieves a performance gain of over 80% in RMSE reduction relative to RF-ISAC, highlighting its robustness and superior estimation capability in noisy environments. As shown in Fig. 5, it is evident that F-MAC consistently achieves a higher communication rate, approximately 304.8 Mbps, regardless of the SNR, whereas RF-ISAC maintains a rate of around 301.2 Mbps. The enhanced communication performance can be attributed to the incorporation of vision-RF multimodal fusion, which enables F-MAC to more accurately locate targets, identify their motion state, and perform more accurate ISAC beam adjustments.
These results validate the effectiveness of multimodal fusion in F-MAC. On one hand, the inclusion of the vision modality enables F-MAC to leverage richer information for sensing tasks, thereby improving sensing accuracy and robustness compared to single-modality RF-based approaches. On the other hand, based on the accurate sensing results, the base station could perform more efficient ISAC beam adjustment, thus improving the communication rate.
VI Open Issues
Despite the promising capabilities of the multimodal ISAC, several challenges remain open for future research and development.
VI-A Adaptive Multimodal Sensing
In dynamic environments, it is essential to adaptively determine which modalities should be activated at each moment or for each task. This requires intelligent policies capable of balancing sensing performance, power consumption, and network load. Therefore, designing lightweight and context-aware modality selection algorithms, possibly informed by reinforcement learning or probabilistic reasoning, remains an open challenge.
VI-B Dynamic Synesthesia Access
Synesthesia access refers to the ability of agents to selectively access different types of sensing sources depending on the situation. It is important that how to make intelligent and timely decisions about which sensors to access for a given task, taking into account the current environment, task requirements, latency constraints, and the availability of sensing resources. Hence, it is crucial to develop unified frameworks capable of dynamically orchestrating such sensor access decisions across multiple agents.
VI-C Online Learning
To maintain performance in changing environments, the agents deployed on the sensors must be capable of evolving. This accounts for efficient online learning strategies that enable the continual update of AI models with minimal communication and computation overhead. Detailed challenges include how to avoid catastrophic forgetting, preserve stability-plasticity balance, and reduce training latency in edge and collaborative settings.
VI-D Privacy and Security
Ensuring data privacy and security in sensors is a critical concern, particularly when the multimodal data includes sensitive visual, audio, or location information. Future research should explore privacy-preserving techniques, such as federated learning, differential privacy, and homomorphic encryption, adapted for multimodal and distributed architectures. Moreover, robust authentication and trust mechanisms are needed to prevent adversarial access or manipulation of sensor data and model outputs.
@@ VII Conclusion
In this article, we first highlighted the necessity of transitioning from conventional single-modal ISAC to multimodal ISAC, driven by the growing demands of complex and dynamic 6G applications. Then, we analyzed the key challenges in enabling multimodal ISAC, including the fusion of multimodal data, the high communication overhead among distributed sensors, and the design of efficient and scalable system architectures. To address these challenges, we introduced several advanced enabling technologies, namely, LAM, SC, and MAS, which collectively provide the foundations for achieving multimodal ISAC schemes. Next, to operationalize these technologies, we zoomed into three architectural paradigms: F-MAC, I-MAC, and R-MAC, each designed to support efficient device coordination and modality collaboration under different deployment scenarios. A case study based on the F-MAC framework demonstrated substantial improvements in sensing accuracy and image reconstruction quality, achieving up to an 80% performance gain compared to conventional RF-ISAC systems. Finally, we outlined several open research challenges that must be addressed in the future.

### v2

@@ V-A Experimental Settings
V-A1 Data Settings
To evaluate the proposed CET framework under realistic multimodal ISAC conditions, we utilize the DeepSense 6G dataset [1], specifically focusing on Scenario 31 (daytime street) and Scenario 33 (nighttime street). These scenarios offer a diverse range of environmental conditions essential for testing robustness. We implement a unified data loading pipeline that supports batched, shuffled, and parallel sampling. To mitigate data imbalance across heterogeneous modalities, mini-batches are sampled from scenario-specific subsets using a probability distribution proportional to the dataset size, ensuring balanced training dynamics. Data preprocessing follows standard protocols in multimodal sensing to guarantee numerical stability: (i) RGB images: Resizing, normalization, and standard augmentation techniques (e.g., random cropping) are applied. (ii) RF power: Sequences undergo min–max normalization and NaN value imputation. (iii) LiDAR point clouds: Data is centered, scaled to a unit sphere, and padded to a fixed point count. (iv) Radar mmWave: Complex-valued IQ signals are normalized and reshaped into consistent sequence formats.
V-A2 Baseline Settings
To validate the efficacy of the proposed CET framework, we benchmark against three distinct operation modes. The primary differentiator is the number of accessible sensing modalities, determined by the mode’s coverage capabilities (as illustrated in Table I).
  • GFM: This represents the full collective mode of cloud-edge-terminal. We assume a centralized cloud server with global coverage and powerful LAMs, enabling it to aggregate and process all available modalities from the terminal: Image (I), RF power (P), Point Cloud (C), and mmWave (M). This setup serves as the theoretical upper bound of sensing accuracy.
  • CRM: This baseline represents the edge-terminal collective mode. We simulate a cooperative scenario involving nearby edge nodes (e.g., RSU or V2V pairs). Through short-range communication, the primary agent can augment its local sensing with an additional modality from a neighbor, expanding the accessible set to three modalities. This mode balances coverage and overhead. The specific variants are:
    1. 1.
      CRM (P+I+C): Fuses RF Power, RGB Image, and LiDAR.
    2. 2.
      CRM (P+I+M): Fuses RF Power, RGB Image, and mmWave.
    3. 3.
      CRM (P+C+M): Fuses RF Power, LiDAR, and mmWave.
  • PIM: This represents the terminal mode. We simulate a standalone resource-constrained device (e.g., NVIDIA Jetson) relying solely on onboard sensors. Due to hardware limitations, it processes only dual-modality inputs. This setting evaluates performance under strict resource constraints. The variants include:
    1. 1.
      PIM (P+I): Fuses RF Power and RGB Images.
    2. 2.
      PIM (P+C): Fuses RF Power and LiDAR Point Clouds.
    3. 3.
      PIM (P+M): Fuses RF Power and mmWave Radar data.
This tiered experimental design aims to quantify the trade-offs between the comprehensive sensing of GFM, the collaborative enhancement of CRM, and the lightweight efficiency of PIM.
V-A3 Metric Settings
We adopt a comprehensive evaluation strategy that balances communication performance with system efficiency. Two primary categories of metrics are considered:
  • Beam Prediction Accuracy: Formulated as a multi-class classification task, this metric evaluates the Top-1 accuracy of the predicted optimal beam index. It serves as a direct proxy for the communication gain (e.g., effective achievable rate) provided by each sensing scheme.
  • Computational Complexity: To assess the deployment feasibility on resource-constrained edge and terminal devices, we analyze the system efficiency through three distinct indicators:
    1. 1.
      Floating Point Operations (FLOPs): This measures the theoretical computational cost of the model, reflecting the processing power required.
    2. 2.
      Memory Footprint: We monitor the peak GPU memory (VRAM) usage during inference to evaluate the spatial complexity and hardware storage requirements.
    3. 3.
      Inference Latency: This records the average end-to-end execution time per sample, which is critical for determining the real-time capability in dynamic ISAC scenarios.
@@ V-B Evaluation Results
V-B1 Performance Analysis
The beam prediction accuracy versus Signal-to-Noise Ratio (SNR) for Scenario 31 (daytime) and Scenario 33 (nighttime) is illustrated in Fig. 4(a) and Fig. 4(b), respectively.
Figure 4: The beam prediction accuracy under different schemes in (a) Scenario 31 and (b) Scenario 33.
It can be observed that the GFM consistently achieves the highest prediction accuracy across all SNR regimes in both scenarios. For instance, in Fig. 4(a), when SNR=25 dB, GFM reaches an accuracy of approximately 76.9%, outperforming all other variants. This superiority is attributed to the GFM’s ability to leverage the comprehensive information from more modalities simultaneously, effectively establishing a performance upper bound for the system. The CRM effectively bridges the performance gap between the cloud and the terminal. As shown in Fig. 4(b), while the performance of the vision-dependent PIM (P+I) degrades significantly due to poor lighting conditions at night, the cooperative CRM (P+I+M) leverages the additional mmWave radar modality from a neighbor to maintain high accuracy, closely approaching the GFM. In contrast, the PIM exhibits varying degrees of limitation. This performance hierarchy (GFM >> CRM >> PIM) highlights the trade-off between modality richness and sensing accuracy.
V-B2 Complexity Analysis
TABLE II: Computational complexity comparison
Method FLOPs (G) Memory (MB) Inference Latency (ms)
PIM(P+I) 4.15e+00 2.40e+01 7.27e+00
PIM(P+C) 9.45e+00 1.05e+02 3.22e+01
PIM(P+M) 7.30e+00 1.89e+01 5.51e+00
CRM(P+I+C) 9.45e+00 1.07e+02 3.11e+01
CRM(P+I+M) 7.31e+00 2.46e+01 1.33e+01
CRM(P+C+M) 8.46e+00 1.06e+02 3.42e+01
GFM 1.26e+01 1.07e+02 3.82e+01
Table II presents the computational complexity per sample in terms of FLOPs, memory footprint, and inference latency (excluding transmission delay).
In terms of resource consumption, the GFM imposes a substantial burden. While this high complexity underpins its superior accuracy, it represents an approximately $3\times$ increase in computation compared to the lightweight PIM (P+I). Such heavy resource demands make the GFM challenging to deploy directly on power-constrained edge devices. Regarding execution speed, the PIM variants demonstrate significant advantages, operating with extremely low inference latency. Notably, the CRM occupies a strategic “sweet spot.” This indicates that CRM can support sophisticated multi-modal fusion without the heavy latency penalty of the full cloud model. Moreover, the practical deployment latency extends beyond computation to include data transmission overhead:
  • GFM: Necessitates uploading the feature data from the terminal to the cloud via edge nodes. Assuming a moderate uplink bandwidth of 50 Mbps, the transmission latency could exceed 300 ms, making the total system latency (>340>340 ms) unsuitable for delay-sensitive beam tracking.
  • CRM: Relies on high-bandwidth, short-range D2D/V2X links (e.g., neighbors). The transmission latency is significantly lower than the cloud uplink, ensuring real-time responsiveness.
  • PIM: Processes data locally, offering the lowest possible total latency ($\sim$ 5-10 ms).
The experimental results validate the motivation of our proposed CET framework. While GFM provides the accuracy upper bound and PIM ensures the latency lower bound, the CRM offers a flexible balance. This necessitates a collaborative architecture that dynamically switches between these modes to satisfy the conflicting requirements of high-precision and low-latency ISAC.
VI Open Issues and Future Directions
In the proposed CET framework, although the integration of LAMs, SC, and MAS enhances sensing efficiency through distributed intelligence, expands the attack surface across the cloud, edge, and terminal layers. These new security challenges transcend traditional physical-layer jamming.
VI-A Semantic Tampering in Global Fusion
In the GFM, terminal agents compress raw data into semantic features for uplink transmission to the cloud. Unlike traditional bit-flipping attacks, adversaries may launch semantic tampering attacks by perturbing the feature vectors in the latent space. Such attacks can impact high-level interpretations (e.g., changing a “pedestrian” label to “background”) without disrupting the communication link or triggering conventional error correction mechanisms, thereby poisoning the global environmental model at the cloud. Therefore, we should explore new semantic integrity verification mechanisms, such as embedding fragile watermarks or cryptographic signatures directly into the semantic latent space, alongside training robust SC models via adversarial learning to resist feature perturbations.
VI-B Malicious Relaying and Instruction Injection in Cooperative Edge
The CRM relies on edge nodes to act as semantic relays, translating observations into control directives for other terminals (e.g., guiding a radar beam based on visual inputs). This may make the system vulnerable to malicious relaying and instruction injection. A compromised edge node could alter the semantic intent of a message or inject fake guidance commands, effectively blinding the sensing system or forcing terminals to waste resources scanning space. To mitigate this, a zero-trust architecture is required where terminal agents cryptographically verify the legitimacy of edge directives. Additionally, behavior-based anomaly detection can be deployed to identify and isolate edge nodes that issue statistically improbable or contradictory control commands.
VI-C Cross-Modal Misleading in Peer Interaction
In the PIM, autonomous terminals exchange high-level semantic alerts directly to facilitate rapid local collaboration. However, this decentralization is vulnerable to cross-modal misleading, where a compromised sensor (e.g., a camera) broadcasts falsified object detections to its peers. Without a central authority to verify ground truth, honest peer agents (e.g., radars) may be tricked into resolving cognitive conflicts in favor of the false data, leading to collective hallucinations. As a solution, we can develop physics-aware consistency checks, where agents utilize physical laws (e.g., correlating Doppler shifts with visual optical flow) to validate peer data. Furthermore, a decentralized reputation management system can be established to dynamically reduce the trust weight of nodes that frequently report uncorroborated semantics.
@@ VII Conclusion
In this article, we argued that multimodal ISAC is a necessary evolution beyond RF-only designs to meet the sensing and connectivity requirements of emerging 6G applications. After summarizing key challenges (heterogeneous fusion, communication overhead, and scalable architectures), we reviewed enabling technologies (LAM, SC, MAS) and proposed a CET framework with three adaptive modes (GFM/CRM/PIM). A case study demonstrated the unique advantages of different modes in this framework, and we outlined open issues for future research.